Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

1. Section 15.1: Problems 7, 19.

2. Section 15.2: Problems 11, 19, 27, 29, 35, 43, 47, 69.

3. Section 15.3: Problems 13, 17.

4. Section 15.4: Problems 11, 21, 23, 26, 41, 42, 43.